A ball is released from a window that is 10 meters above the ground. When the ball leaves your hand, it is moving 8m/s at an angle 300 below the horizontal. Ignoring the effect of air resistance, how far horizontally from the window will the ball hit the ground?


1
Expert's answer
2021-08-24T02:54:37-0400

"U_x=ucos\u03b1=8*cos30=6.92820m\/s\\\\ U_y=usin\u03b1=8*sin30 =4m\/s\\\\ Y=usin\u03b1*t+\\frac{1}{2}gt^2\\\\ 10=4t+6.92820t^2\\\\ Solving \\space for \\space t=0.95sec\\\\ Horizontal \\space distance=U_x*t=6.92820*0.95=6.58179mU \nx\n\u200b\n =ucos\u03b1=8\u2217cos30=6.92820m\/s\nU \ny\n\u200b\n =usin\u03b1=8\u2217sin30=4m\/s\nY=usin\u03b1\u2217t+ \n2\n1\n\u200b\n gt \n2\n \n10=4t+6.92820t \n2\n \nSolving for t=0.95sec\nHorizontal distance=U \nx\n\u200b\n \u2217t=6.92820\u22170.95=6.58179m"